OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n. 10 + 30 + 60 + ... + 10n = 5n(n + 1)

OpenStudy (anonymous):

Call the statement 10 + 30 + 60 + ... + 10n = 5n(n + 1) P(n), and show using induction the statement is true for all positive integers n. Basis: show the statement holds true for n=0 10 (0) = 5*0(0+1) 0 = 0 The basis holds for P(0) Inductive Hypothesis: If P(k) holds for some unspecified value of k, then so does P(k+1); it must be shown that (10 + 30 + 60 ..... + 10k) + 10(k+1) = 5(k+1)((k+1)+1) Using the inductive hypothesis that P(k+1) holds, we can say that the left hand can be written as 5k(k+1) + 10(k+1) = 5k^2 + 5k + 10k +10 = 5k^2 + 10k + 5 + 5k + 5 = 5(k^2 + 2k +1 + k + 1) = 5((k^2+2k +1)) +(k+1)) = 5((k + 1)(k + 1) + (k + 1)) = 5(k+1)((k+1) + 1) Therefor P(k+1) holds algebraically, and since the basis and inductive step have been performed, the statement P(n) holds for all positive integers.